OpenStudy (anonymous):

medal!!!!!!!!!!!!!!!!!!!!!!!! Amy can drive 22.8 miles on one gallon of gas. How far can Amy drive on 5.3 gallons of gas?

OpenStudy (anonymous):

120.84

OpenStudy (anonymous):

?

OpenStudy (paki):

use ratio to solve... 22.38 : 1 x : 5.3 solve for "x" @isabell2003

OpenStudy (hanner_b_nanner):

yes, you're right! @isabell2003

OpenStudy (texaschic101):

22.8/1 = x/5.3 -- (22.8 miles to 1 gallon = x miles to 5.3 gallons) cross multiply x = (22.8)(5.3) x = 120.84 yep...your correct